# Deep Unfolding Network with Spatial Alignment for multi-modal MRI reconstruction

### v1

@@ 7 Conclusion
In this paper, we propose a novel joint alignment and reconstruction model for multi-modal MRI reconstruction. By developing a cross-modal spatial alignment term, we integrate the spatial alignment task into the reconstruction process. We design an optimization algorithm for solving it and then unfold each iterative step into the corresponding network module. As a result, we have constructed a deep unfolding network with interpretability, termed DUN-SA. Through end-to-end training, we fully leverage both intra-modality and inter-modality priors. Comprehensive experiments conducted on three real datasets have demonstrated that the proposed DUN-SA outperforms current state-of-the-art methods in both quantitative and qualitative assessments. Additionally, we have verified that DUN-SA is relatively robust to misalignment, with minimal impact on spatial alignment even as acceleration factors increase.

### v2

@@ 7 Conclusion
In this paper, we propose a novel joint alignment and reconstruction model for multi-modal MRI reconstruction. By developing an aligned cross-modal prior term, we integrate the spatial alignment task into the reconstruction process. We design an optimization algorithm for solving it and then unfold each iterative stage into the corresponding network module. As a result, we have constructed a deep unfolding network with interpretability, termed DUN-SA. Through end-to-end training, we fully leverage both intra-modality and inter-modality priors. Comprehensive experiments conducted on four real datasets have demonstrated that the proposed DUN-SA outperforms current state-of-the-art methods in both quantitative and qualitative assessments. Additionally, we have verified that DUN-SA is relatively robust to misalignment, with minimal impact on spatial alignment even as acceleration factors increase.